Simplifying the expression 3/4+1/2×1/6 gives?




1
Expert's answer
2021-04-28T08:49:00-0400

"\frac{3}{4}+\frac{1}{2}*\frac{1}{6}=\frac{3}{4}+\frac{1}{12}=\frac{9}{12}+\frac{1}{12}=\frac{10}{12}=\frac{5}{6}" .